OpenStudy (anonymous):

integral of cos^3xcos(2x)dx

OpenStudy (kira_yamato):

\[\cos3x = =4 \cos^3 x - 3\cos x \] Use integration by parts

OpenStudy (anonymous):

how did you arrive to 4cos^3x

OpenStudy (kira_yamato):

Triple angle formula for cosines is given in my last reply. Add 3cosx on each side and then divide both sides of the equation (or identity) to get the following: \[\cos^3x = \frac{\cos3x + 3\cos x}{4}\]

OpenStudy (kira_yamato):

If you want to derive the triple angle formula, use cos 3x = cos (x + 2x), then the addition formula, and finally the double angle formula...

OpenStudy (anonymous):

Triple angle formula? I don't think we cover that in my calculus 2 class... but I knew it's suppose to be solve with integral by parts

OpenStudy (kira_yamato):

Triple angle formula is under trigonometry...

OpenStudy (anonymous):

Thanks for the help. I have another question. \[\int\limits_{0}^{\infty} (16\tan^-1)/1+x^2\]. I know it's solve by using trig substitution but how do you set it up. The Tan inverse is throwing me off

OpenStudy (kira_yamato):

Sorry... I don't know either...